OpenStudy (anonymous):

whats the GCF of 72,48 an 36?

OpenStudy (anonymous):

Do you know the euclidean algorithm?

OpenStudy (anonymous):

You can also just decompose the numbers into prime factors. Choose either method.

OpenStudy (anonymous):

\[72 = 2^3 \cdot 3^2\]\[48 = 2^4 \cdot 3\]\[36 = 2^2 \cdot 3^2\] We take the greatest powers of the prime factors common to all the numbers: \[GCF = 2^2 \cdot 3 = 12\] Let me know if you have any problems understanding the answer.